On the road map with a scales of 1/4 inch per 10 miles, the highway from Willkes baree to Scranton is 1-3/8 inches long. How many miles long is this highway?
Answer by rfer(16322)

2/8=10
1/8=5
distance is 11/8
5*(11)=55 mi